find the equation for the line passing through the point (1,-4) and parallel to the line whose equation is y=4x+10
Answer by Tatiana_Stebko(1539)

An equation of a line through a point P(x1 , y1) with slope m is given by
y-y1=m(x-x1)
Point (1,-4), then x1=1 and y1=-4
If the line is parallel to the line whose equation is y=4x+10, then their slopes are equal, so m=4
y-(-4)=4(x-1)
y+4=4x-4
y=4x-8 the equation of a line